Reports 1-1 of 1 Clear search Modify search
DetChar (General)
takaaki.yokozawa - 8:42 Thursday 29 August 2019 (10182) Print this report
Noise investigation 2019/08/29 FPMI


- Not serious coherence at 7.5Hz, appear the coherence at the ETMY_TM_DAMP_P_IN(No TM P damp) and ETMY oplev pitch and MICH control singal (very small in SR2 L damp)

- 10-20Hz Type-A part limit?? ->See figure1

- current limit 10-  by VIS-MCE_TM_SICINF_L -> See figure2

- 20-50 Hz MICH controls is also large coherence

- 40.5Hz large peak ITMX Yaw??

- 46.6Hz no clear coherence was detected from ITMX pitck

Images attached to this report
Search Help